OpenStudy (anonymous):

find an equation for the nth term of the geometric sequence where the second and fifth terms are -8 and 512, respectively. A. an=2*(-4)^n+1 B. an=2*4^n-1 C. an=2*(-4)^n-1 D. an=2*4^n @dan815 @michele_laino @nincompoop @campbell @paki @nnesha @e.mccormick

OpenStudy (anonymous):

@ash2326 @welshfella @KyanTheDoodle @mathmath333 @misssunshinexxoxo

OpenStudy (welshfella):

if a1 is the first term and r = common ratio second term = a1r = -8 fifth term = a1r^4 = 512 dividing a1r^4 / a1r = 512/-8 r^3 = -64 therefor r = -4

OpenStudy (welshfella):

now you can see that a1*-4 = -8 so now you can find what a1 is

OpenStudy (anonymous):

would it be c?

OpenStudy (welshfella):

nth term = a1r^(n-1) so an = 2*(-4)^(n-1)

OpenStudy (welshfella):

yea c is correct

OpenStudy (anonymous):

can u help me with another one?